# Globalization of perturbative Chern-Simons theory on the moduli space of flat connections in the BV formalism

### v2

@@ 2.1.2. Lifting to forms
We have the following result.
Proposition 2.6.
The formal power series $\widetilde{\delta}_{A_{0}}$ defines a Maurer-Cartan element in $\Omega^{1}[[t]]$:
(51) $d_{A_{0}}\widetilde{\delta}_{A_{0}}+\frac{1}{2}[\widetilde{\delta}_{A_{0}},\widetilde{\delta}_{A_{0}}]=0.$
We will prove equation (51) using smoothness of $\mathrm{FC}$ at $A_{0}$ which we now establish by defining another lift of $\alpha$ to a formal deformation. Namely, we split a closed 1-form $\alpha$ as $\alpha=i_{A_{0}}p_{A_{0}}\alpha+d_{A_{0}}K_{A_{0}}\alpha$. Denoting $\beta=K_{A_{0}}\alpha$, we can lift $\alpha$ to a (formal) curve of flat connections with tangent vector $\dot{A}_{0}=\alpha$ by setting $\mathrm{g}_{t}=\exp(-t\beta)$ and setting
@@
with $\alpha_{h}^{(l)}$ as in Proposition 2.4 applied to $\mathsf{a}=p_{A_{0}}\alpha$. In this way, we have constructed the map $\widetilde{\psi}_{A_{0}}(t\alpha)\colon\Omega^{1}\to\Omega^{1}[[t]]$ lifting any infinitesimal deformation of $A_{0}$, i.e. a closed form, to a formal deformation.
@@ 2.1.3. Convergence in Banach norm
Suppose $K_{A_{0}}$ is bounded with respect to a Banach norm $||\cdot||_{B}$ on $\Omega^{\bullet}$ (Assumption 2.2). Then there is an open interval II around zero such that $\widetilde{\delta}_{A_{0}}$ defined by (50), seen as a formal power series in $t$, converges in $||\cdot||_{B}$ for $t\in I$.
@@
The summand $\alpha^{(n)}$ is a sum over binary trees with $n$ leaves, involving at most $n$ applications of $K_{A_{0}}$. The number of such binary trees is given the (n1)(n-1)-th Catalan number, and by assumption $K_{A_{0}}$ is bounded in $||\cdot||_{B}$, say with constant $L$. The lie bracket is bounded by the maximum of the structure constants, say $F$. We therefore have the simple estimate
@@
Equivalently, we can set $t=1$ by letting $\alpha||_{B}$ small enough, i.e. there exists some open set $\widetilde{U}_{A_{0}}\subset\Omega^{1}$ on which the formal power series $\widetilde{\delta}_{A_{0}}(\alpha)$ converges, i.e. there is a map
(54) $\widetilde{\delta}_{A_{0}}\colon\Omega^{1}\supset\widetilde{U}_{A_{0}}\to\overline{\Omega^{1}}$
given by (50) with $t=1$. A priori, the limit of this power series lives in the completion $\overline{\Omega^{1}(M,\mathfrak{g})}$ of $\Omega^{1}(M,\mathfrak{g})$ with respect to $||\cdot||_{B}$. In the case of interest to use in the paper, however, the limit is smooth by nonlinear elliptic regularity. Convergence of the map $\widetilde{\delta}_{A_{0}}$ implies that there are well-defined maps
@@ 2.1.4. Kuranishi map
(59) $\widetilde{\kappa}_{A_{0}}(\delta)=\delta+\frac{1}{2}K_{A_{0}}[\delta,\delta].$
Some salient properties of the this map are:
@@
Assume $r_{A_{0}}$ satisfies the Banach boundedness assumption 2.2. Then:
@@
    The map $\widetilde{\kappa}_{A_{0}}\colon\Omega^{1}_{B}\to\Omega^{1}_{B}$ is invertible in a neighborhood $V_{A_{0}}$ of $0\in\Omega^{1}_{B}$, and V_A_0 ⊃{δΩ^1—  ——K_A_0ad_δ——_op ¡ 1}.
@@
The proof is in Appendix D.1
Lemma 2.10.
@@
We know that $\widetilde{\delta}_{A_{0}}$ converges in a neighborhood $U_{1}$ of $0\in\Omega^{1}_{B}$, therefore $\widetilde{\delta}_{A_{0}}$ is defined on $U=U_{1}\cap\operatorname{Im}i_{A_{0}}$. On $\operatorname{Im}i_{A_{0}}$, we know that $\widetilde{\delta}_{A_{0}}\in MC_{A_{0}}$ by Proposition 2.4. Since $\widetilde{\delta}_{A_{0}}(\alpha)=\alpha+K_{A_{0}}(\ldots)$, we have $K_{A_{0}}\widetilde{\delta}_{A_{0}}(\alpha)=0$. Therefore $\widetilde{\delta}_{A_{0}}(U)\subset MC_{A_{0}}\cap\ker K_{A_{0}}$. On the other hand, if $m\in MC_{A_{0}}\cap\ker K_{A_{0}}$, then $d_{A_{0}}\widetilde{\kappa}(m)=K_{A_{0}}\widetilde{\kappa}(m)=0$, i.e $\widetilde{\kappa}_{A_{0}}(m)\in\operatorname{Im}i_{A_{0}}$. For $m$ small enough, we therefore have $\widetilde{\kappa}_{A_{0}}m\in U$, and since we already know that $\widetilde{\delta}_{A_{0}}$ and $\widetilde{\kappa}_{A_{0}}$ are inverse to each other, we conclude the statement. ∎
@@
The restriction of the map $\kappa_{A_{0}}\colon{\Omega^{1}(M,g)}\to H^{1}_{A_{0}}(M,\mathfrak{g})$ given by
(61) $\kappa_{A_{0}}(\delta)=p_{A_{0}}(\delta+\frac{1}{2}K_{A_{0}}[\delta,\delta])=p_{A_{0}}(\delta)\in H^{1}_{A_{0}}(M,\mathfrak{g})$
We know that $\widetilde{\psi}_{A_{0}}(\alpha)$ is flat for any closed 1-form $\alpha$ in its domain. We claim that locally, it is actually invertible and thus provides $FC$ with the structure of a Banach manifold around $A_{0}$.
Proposition 2.12.
There exists a neighborhood $U_{A_{0}}$ of 0 in $\Omega^{1}_{A_{0}-\mathrm{cl},B}$ such that the map $\widetilde{\psi}_{A_{0}}\colon U_{A_{0}}\to\mathrm{FC}_{B}\subset\Omega^{1}_{B}$ is a homeomorphism onto its image.
We give the proof below, but first record that together with the above discussion, we have the following corollary:
@@
The subset of smooth points $\mathrm{FC}^{\mathrm{sm}}_{B}$ is a Banach manifold. For each point $A_{0}\in\mathrm{FC}^{\mathrm{sm}}_{B}$, there is a local chart $\widetilde{\varphi}_{A_{0}}\colon\Omega^{1}_{A_{0}-\mathrm{cl}}\supset U_{A_{0}}\to\mathrm{FC}_{B}$.
In particular, Theorem 2.3 follows.
Proof of Proposition 2.12.
To show surjectivity onto a small neighborhood of $A_{0}\in\mathrm{FC}_{B}$, in the first step, we construct a gauge transformation that takes an arbitrary flat $A_{1}$ close enough to $A_{0}$ to a connection $A_{1}^{\prime}$ satisfying $K_{A_{0}}(A_{1}^{\prime}-A_{0})=0$. To this end, consider the map $F\colon\Omega^{0}_{K-\mathrm{ex}}\times\Omega^{1}(M,\mathfrak{g})\to\Omega^{0}_{K-\mathrm{ex}}$ given by
$(\beta,\delta)\to K_{A_{0}}(\,{}^{\exp(-\beta)}(A_{0}+\delta)-A_{0}).$
We want to solve for $\beta=\beta(\delta)$ such that $F(\beta,\beta(\delta))=0$, this is the desired gauge transformation. Existence of $\beta$, for small enough $\delta$, is then guaranteed by the implicit function theorem for Banach spaces, since the derivative of $F$ at $(0,0)$ in direction of $\beta$ is $(dF/d\beta)(0,0)=K_{A_{0}}d_{A_{0}}=\mathrm{id}_{\Omega^{0}_{K-\mathrm{ex}}}$. This means that for $A_{1}$ close enough to $A_{0}$, there is $\beta$ such that $K_{A_{0}}(\,{}^{\exp(-\beta)}A_{1}-A_{0})=0$. For such connections, we know that $\widetilde{\kappa}_{A_{0}}(\,{}^{\exp(-\beta)}A_{1}-A_{0})$ is a $d_{A_{0}}$- and $K_{A_{0}}$-closed 1-form. Therefore, if $A_{1}$ is a flat connection close to $A_{0}$, then $\alpha=\widetilde{\kappa}_{A_{0}}(\,{}^{\exp(-\beta)}A_{1}-A_{0})-d_{A_{0}}\beta$ is a $d_{A_{0}}$-closed form such that $\widetilde{\psi}_{A_{0}}(\alpha)=A_{1}$. ∎
@@ 2.2. Transporting “harmonic” forms
Suppose we are given a smooth flat connection $A_{0}$ and a harmonic 1-form $\mathsf{a}$. Let $A_{t}=A_{0}+\widetilde{\delta}_{A_{0}}(t\mathsf{a})=\widetilde{\varphi}_{A_{0}}(t\mathsf{a})$ be the path of flat connections given by Proposition 2.4. Let $\chi$ be another harmonic form. Then, there exists a deformation
@@
2.3. Irreducible points
Definition 2.18.
Example 2.19.
@@ 2.5.1. Good gauge fixing operators
The main example of a good gauge fixing operator is, given the choice of a Riemannian metric on $M$, the codifferential $d^{*}_{A_{0}}$. The fact that $d^{*}_{A_{0}}$ is a good gauge-fixing operator follows from the well-known Hodge decomposition. In this case

### v3

@@ 2.1.3. Lifting to forms
We now establish by defining another lift of $\alpha$ to a formal deformation. Namely, we split a closed 1-form $\alpha$ as $\alpha=i_{A_{0}}p_{A_{0}}\alpha+d_{A_{0}}K_{A_{0}}\alpha$. Denoting $\beta=K_{A_{0}}\alpha$, we can lift $\alpha$ to a (formal) curve of flat connections with tangent vector $\dot{A}_{0}=\alpha$ by setting $\mathrm{g}_{t}=\exp(-t\beta)$ and setting
@@
with $\alpha_{h}^{(l)}$ as in Proposition 2.7 applied to $\mathsf{a}=p_{A_{0}}\alpha$. In this way, we have constructed the map $\widetilde{\psi}_{A_{0}}(t\alpha)\colon\Omega^{1}_{A_{0}-\mathrm{cl}}\to\Omega^{1}[[t]]$ lifting any infinitesimal deformation of $A_{0}$, i.e. a closed form, to a formal deformation.
@@ 2.1.4. Convergence in Banach norm
Suppose $K_{A_{0}}$ is bounded with respect to $||\cdot||_{l}$ for some $l\geq 2$ on $\Omega^{\bullet}$ (Assumption 2.3). Then there is an open interval II around zero such that $\widetilde{\delta}_{A_{0}}$ defined by (51), seen as a formal power series in $t$, converges in $||\cdot||_{l}$ for $t\in I$.
@@
The summand $\alpha^{(n)}$ is a sum over binary trees with $n$ leaves, involving at most $n$ applications of $K_{A_{0}}$. The number of such binary trees is given the (n1)(n-1)-th Catalan number, and by assumption $K_{A_{0}}$ is bounded in $||\cdot||_{l}$, say with constant $L$. The Lie bracket $\alpha,\alpha]$ is bounded by
@@
[…]
Equivalently, we can set $t=1$ by letting $\alpha||_{l}$ small enough, i.e. there exists some open set $\widetilde{U}_{A_{0}}\subset\Omega^{1}$ on which the formal power series $\widetilde{\delta}_{A_{0}}(\alpha)$ converges, i.e. there is a map
(55) $\widetilde{\delta}_{A_{0}}\colon\Omega^{1}_{l}\supset\widetilde{U}_{A_{0}}\to\Omega^{1}_{l}$
given by (51) with $t=1$.
Remark 2.10.
A priori, the limit of this power series lives in the completion $\Omega^{1}_{l}$ of $\Omega^{1}$ with respect to $||\cdot||_{l}$. However, under the additional Assumption 2.4 of Fréchet continuity the sum actually converges in the Fréchet topology and $\widetilde{\delta}_{A_{0}}$ actually lands in $\Omega^{1}$.
Convergence of the map $\widetilde{\delta}_{A_{0}}$ implies that there are well-defined maps
@@ 2.1.5. Kuranishi map
(60) $\widetilde{\kappa}_{A_{0}}(\mathfrak{d})=\mathfrak{d}+\frac{1}{2}K_{A_{0}}[\mathfrak{d},\mathfrak{d}].$
This map is as regular as $K_{A_{0}}$, in particular if $K_{A_{0}}$ is $l$-Sobolev bounded (for $l\geq 2$) then so is $\widetilde{\kappa}_{A_{0}}$ and if $K_{A_{0}}$ is Fréchet continuous then so is $\widetilde{\kappa}_{A_{0}}$. Some salient properties of the this map are:
@@
Assume $r_{A_{0}}$ satisfies the $l$-Sobolev boundedness assumption for some $l\geq 2$ (Assumption 2.3). Then $\widetilde{\kappa}_{A_{0}}$ extends to the completion $\Omega^{1}_{l}$ and we have:
@@
    The map $\widetilde{\kappa}_{A_{0}}\colon\Omega^{1}_{l}\to\Omega^{1}_{l}$ is invertible in a neighborhood $V_{A_{0}}$ of $0\in\Omega^{1}_{l}$, and
[…]
    Here $||\cdot||_{\mathrm{op}}$ denotes the operator norm induced by the Sobolev norm $||\cdot||_{l}$.
@@
  4. iv)
    We have
    (62) $\widetilde{\varphi}_{A_{0}}^{-1}(A_{1})=\widetilde{\kappa}_{A_{0}}(A_{1}-A_{0}).$
The proof is in Appendix D.1.1.
2.1.6. Banach and Fréchet smoothness
We know that $\widetilde{\psi}_{A_{0}}(\alpha)$ is flat for any closed 1-form $\alpha$ in its domain. We claim that locally, it is actually invertible and thus provides $\mathrm{FC}$ with the structure of a Banach manifold around $A_{0}$.
Proposition 2.13.
  1. i)
    There exists a neighborhood $U_{A_{0}}$ of 0 in $\Omega^{1}_{A_{0}-\mathrm{cl},l}$ such that the map $\widetilde{\psi}_{A_{0}}\colon U_{A_{0}}\to\mathrm{FC}_{l}\subset\Omega^{1}_{l}$ is a homeomorphism onto its image.
  2. ii)
    If $A_{1}\in\widetilde{\psi}_{A_{0}}(U_{A_{0}})$, then the composition $\widetilde{\psi}_{A_{1}}^{-1}\circ\widetilde{\psi}_{A_{0}}$ is smooth [^35] In the Banach space setting this means smoothness w.r.t. the Fréchet derivative. on its domain.
In particular, $\mathrm{FC}^{\mathrm{sm}}_{l}$ is a Banach manifold and $\{\widetilde{\psi}_{A_{0}}\}_{A_{0}\in\mathrm{FC}^{\mathrm{sm}}_{l}}$ is an atlas for $\mathrm{FC}^{\mathrm{sm}}_{l}$. We give the proof in Appendix D.1.2. We then have the following corollary.
Proposition 2.14.
There is a neighborhood $0\in V_{A_{0}}\subset\Omega^{1}_{A_{0}-\mathrm{cl},l}$ such that $\widetilde{\delta}_{A_{0}}$ applied to a closed 1-form $\alpha\in V_{A_{0}}$ defines a Maurer-Cartan element in $\Omega^{1}_{A_{0},l}$:
(63) $d_{A_{0}}\widetilde{\delta}_{A_{0}}(\alpha)+\frac{1}{2}[\widetilde{\delta}_{A_{0}}(\alpha),\widetilde{\delta}_{A_{0}}(\alpha)]=0.$
@@ 2.1.6. Banach and Fréchet smoothness
We know that $\mathrm{FC}^{\mathrm{sm}}_{l}$ is a Banach manifold and for any $A_{0}\in\mathrm{FC}^{\mathrm{sm}}$, the map $A_{1}\mapsto\widetilde{\kappa}_{A_{0}}(A_{1}-A_{0})$ sends $A_{1}\in\mathrm{FC}^{\mathrm{sm}}_{l}$ to $\Omega^{1}_{A_{0}-\mathrm{cl},l}$ (point iii of Proposition 2.12). We can now invoke the inverse function theorem to conclude that this map is surjective onto a neighborhood $V_{A_{0}}$ of zero. We then have for all $\alpha\in V_{A_{0}}$ that $\alpha=\widetilde{\kappa}_{A_{0}}(m)$, with $m\in MC_{A_{0}}$. Therefore, $\widetilde{\delta}_{A_{0}}(\alpha)=\widetilde{\delta}_{A_{0}}\widetilde{\kappa}_{A_{0}}(m)=m$ satisfies the Maurer-Cartan equation. ∎
This implies the following corollary.
@@
For each point $A_{0}\in\mathrm{FC}^{\mathrm{sm}}_{l}$, the map $\widetilde{\varphi}_{A_{0}}\colon\Omega^{1}_{A_{0}-\mathrm{cl}}\supset U_{A_{0}}\to\mathrm{FC}_{l}^{\mathrm{sm}}$ is a chart with inverse given by Equation (62).
In particular, Theorem 2.5 follows. Finally, if $K_{A_{0}}$ is Fréchet continuous, we also have smoothness of $\mathrm{FC}^{\mathrm{sm}}$ in the Fréchet topology:
@@
The set of smooth points $\mathrm{FC}^{\mathrm{sm}}$ is a Fréchet manifold. If $K_{A_{0}}$ satisfies Assumption 2.4, then for each point $A_{0}\in\mathrm{FC}^{\mathrm{sm}}$, restriction of $\widetilde{\varphi}_{A_{0}}$ to $\mathrm{FC}^{\mathrm{sm}}$ defines a chart $\widetilde{\varphi}_{A_{0}}\colon\Omega^{1}_{A_{0}-\mathrm{cl}}\supset U_{A_{0}}\to\mathrm{FC}^{\mathrm{sm}}$.
Proof.
The maps $\widetilde{\kappa}_{A_{0}}$ and $\widetilde{\delta}_{A_{0}}$ are both Fréchet continuous and inverse to each other, hence they are inverse homeomorhisms in the Fréchet topology when restricted to $\mathrm{FC}^{\mathrm{sm}}$ and $\Omega^{1}_{A_{0}-\mathrm{cl}}$ respectively. Also, note that $\widetilde{\kappa}_{A_{0}}$ and $\widetilde{\delta}_{A_{1}}$ are both smooth when considered on open neighborhoods of zero in $\Omega^{1}$. Therefore, if $A_{1}\in\widetilde{\varphi}_{A_{0}}(U_{A_{0}})$, then the composition $\widetilde{\varphi}^{-1}_{A_{1}}\circ\widetilde{\varphi}_{A_{0}}$ is smooth as a map between neighborhoods of zero in $\Omega^{1}$. [^36] In Fréchet spaces derivatives are defined using directional (Gateaux) derivatives for which the chain rule holds. Therefore, its restriction to the subspace $\Omega^{1}_{A_{0}-\mathrm{cl}}$ is also smooth. ∎
This finishes the proof of Theorem 2.6.
@@ 2.3. Transporting “harmonic” forms
Suppose we are given a smooth flat connection $A_{0}$ and a harmonic 1-form $\mathsf{a}$. Let
(65) $A_{t}=A_{0}+\widetilde{\delta}_{A_{0}}(t\mathsf{a})=\widetilde{\varphi}_{A_{0}}(t\mathsf{a})$
be the path of flat connections given by Proposition 2.7. Let $\chi$ be another harmonic form. Then, there exists a deformation
@@
Under the additional assumption that $A_{0}$ is irreducible, Proposition 2.20 can be generalized to paths $A_{t}$ not restricted to the form (65):
Proposition 2.23.
Let $A_{0}\in\mathrm{FC}^{\prime}$ be a smooth irreducible flat connection and let $A_{t}=A_{0}+\alpha_{t}$ with $\alpha_{t}$ be a path of flat connections with $\alpha_{0}=0$. Fix a harmonic form $\chi$. Then the path of forms
(70) $\chi_{t}=\sum_{k\geq 0}(-K_{A_{0}}\mathrm{ad}_{\alpha_{t}})^{k}\circ\chi$
starting at $\chi_{0}=\chi$ satisfies $d_{A_{t}}\chi_{t}=K_{A_{0}}\chi_{t}=0$.
Proof.
Property $K_{A_{0}}\chi_{t}=0$ is immediate by construction (70). The nontrivial part is proving $d_{A_{t}}\chi_{t}=0$.
Due to irreducibility of $A_{0}$, we only have harmonic 1-forms and 2-forms. We consider these two cases for $\chi$ separately.
Let $\chi$ be a harmonic 1-form. Then $\chi_{t}=(d\widetilde{\varphi}_{A_{0}})_{\mathsf{a}_{t}}(\chi)$, where $\mathsf{a}_{t}$ is a path of closed forms such that $\widetilde{\varphi}_{A_{0}}(\mathsf{a}_{t})=\alpha_{t}$, cf. Theorem 2.5. Since $\widetilde{\varphi}_{A_{0}}(\alpha)$ maps closed 1-forms to flat connections, its derivative maps closed 1-forms to first order deformations of flat connections, i.e., $d_{\widetilde{\varphi}_{A_{0}}(\alpha)}$-closed forms. Therefore, $d_{A_{t}}\chi_{t}=0$.
Now, let $\chi$ be a harmonic 2-form. We have
$d_{A_{0}}\chi_{t}=\underbrace{d_{A_{0}}\chi}_{0}-\sum_{k\geq 1}d_{A_{0}}K_{A_{0}}\mathrm{ad}_{\alpha_{t}}(-K_{A_{0}}\mathrm{ad}_{\alpha_{t}})^{k-1}\circ\chi\\ =-\sum_{k\geq 1}\mathrm{ad}_{\alpha_{t}}(-K_{A_{0}}\mathrm{ad}_{\alpha_{t}})^{k-1}\circ\chi=-\mathrm{ad}_{\alpha_{t}}\chi_{t},$
which implies $d_{A_{t}}\chi_{t}=0$. Here in transition to the second line, we note that the operator $d_{A_{0}}K_{A_{0}}=\mathrm{id}-i_{A_{0}}p_{A_{0}}-K_{A_{0}}d_{A_{0}}$ acts on a 3-form $\beta$, and thus this operator acts as identity ($p_{A_{0}}\beta=0$ by irreducibility of $A_{0}$ and $d_{A_{0}}\beta=0$ by degree reason). ∎
@@ 2.5.1. Good gauge fixing operators
The main example of a good gauge fixing operator is, given the choice of a Riemannian metric on $M$, the codifferential $d^{*}_{A_{0}}$. The fact that $d^{*}_{A_{0}}$ is a good gauge-fixing operator follows from the well-known Hodge decomposition. [^40] See e.g. [RS71, Section 3], [MW11] for details on Hodge-de Rham decomposition theorem for forms on a Riemannian manifold with coefficients in a flat bundle. In this case
@@
Proposition 2.31.
On a compact manifold $M$, for any $A_{0}$ and $k\in\mathbb{N},p\geq 2$, the SDR data $(i_{A_{0}},p_{A_{0}},K_{A_{0}})$ is bounded with respect to the Sobolev norm $||\cdot||_{k,p}$. In particular, the SDR data is Fréchet continuous (satisfies Assumption 2.4
Proof.
The maps $i_{A_{0}}$ and $p_{A_{0}}$ are bounded since they have finite-dimensional domain (resp. codomain). Denoting the completion of $\Omega^{\bullet}$ with respect to $||\cdot||_{k,p}$ by $\Omega^{\bullet}_{k,p}$, the map $K$ is bounded since it is a composition of the bounded maps $G_{A_{0}}\colon\Omega^{\bullet}_{k,p}\to\Omega^{\bullet}_{k+2,p}$ and $d^{*}_{A_{0}}\colon\Omega^{\bullet}_{k+2,p}\to\Omega^{\bullet-1}_{k+1,p}$. Therefore it defines a continuous map $K_{A_{0}}\colon\Omega^{\bullet}_{k,p}\to\Omega^{\bullet-1}_{k+1,p}$, and the injection $\Omega^{\bullet-1}_{k+1,p}\to\Omega^{\bullet-1}_{k,p}$ is continuous, which proves the first claim.  [^41] Since the inclusion $\Omega^{\bullet-1}_{k+1,p}\to\Omega^{\bullet-1}_{k,p}$ is compact on a compact manifold by the Rellick-Kondranov theorem, we have actually proven that $K_{A_{0}}$ is compact. Since it is bounded with respect to all Sobolev norms, it is also continuous in the Fréchet topology. ∎